Mavie has two jobs, her permanent job which pays Php 800 per day and her 2-hour part time job which pays her Php200 per day. How many days should she works to earn at least
Php10,000?

Answer by ikleyn(52921)

.

    800d + 200d >= 10000


    1000d       >= 10000


        d       >= 10000/1000 = 10.


ANSWER.  At least 10 days.

Solved.

As simple as 2 + 2 = 4.